OpenStudy (anonymous):

A, B and C are 2x2 matrices a) Given that AB=AC and that A is not singular, prove that B=C b) Given that AB=AC where A=(3 6) and B=(1 5) (1 2) (0 1) find a matrix C whose elements are all non-zero. First bit is fine but surely that means C is B and can't have non zero elements?

OpenStudy (anonymous):

\[B= \left[\begin{matrix}1 & 5 \\0 & 1\end{matrix}\right] A= \left[\begin{matrix}3 & 6 \\ 1 & 2\end{matrix}\right]\]

OpenStudy (beginnersmind):

"First bit is fine but surely that means C is B and can't have non zero elements?" Can you rephrase your question? Are you talking about part a) or part b)? Note that part a) assumes A is non-singular. But in part b) det(A)=3*2-1*6=0, so A is singular.

OpenStudy (anonymous):

Well I was able to get why B=C. But don't worry, if A is singular then C isn't B. How then do I get C from B and A?

OpenStudy (phi):

for part (b), notice that the 2nd col of A is twice the 1st column AB \[\left[\begin{matrix}3 & 6 \\ 1 & 2\end{matrix}\right]\left[\begin{matrix}1 & 5 \\ 0 & 1\end{matrix}\right]\]gives \[\left[\begin{matrix}3 & 21 \\ 1 & 7\end{matrix}\right]\] one way to interpret matrix multiplication is the 1st column of B says that the 1st column in the product matrix will be a linear combination of 1st col of A scaled by 1 and the 2nd col of A scaled by 0. i.e. \[\left[\begin{matrix}3 \\ 1 \end{matrix}\right]\] now, because the 2nd col of A is a scaled version (twice) of the 1st column we could say: use the 2nd col of A and subtract the 1st col of A . This has the effect of 2*1st col - 1*1st col (because 2nd col = 2*1st col) = 1st col A (which is what we want) so one of many answers is \[C=\left[\begin{matrix}-1 & 5 \\ 1 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

Ahh, clever. I get you. Not really technical maths, just realising that 6-3 is the same as 3-0. What other versions of C could there be? Surely thats the only one that works with both 3 and 6 and 1 and 2

OpenStudy (phi):

if we call the 1st col v the 2nd col is 2v what combinations of av+2bv= v ? where a and b are the entries in the 1st col of matrix C solving: a+2b=1 a= 1-2b there are an infinite number of choices: if b =0 a=1, if b=1, a= -1, if b=1/2, a=0...

OpenStudy (anonymous):

Yes, I understand that you can get infinite for 3 and 6 or 1 and 2 but not both surely.

OpenStudy (phi):

If I am not clear: matrix A has column v and column 2v we want the answer matrix C to have v as its first column

OpenStudy (anonymous):

You mean AC to have v as it's first column

OpenStudy (anonymous):

Give me an example of C other than the one you gave where it would give the same matrix as AB

OpenStudy (phi):

yes, I mean AC if you are asking about the 2nd column of AC, yes there are infinite combinations for it also

OpenStudy (anonymous):

No, I'm saying that there are not an infinite number of combinations as you have 6 and 3 and 1 and 2

OpenStudy (anonymous):

So having -3 and 4 in the first column of C would give 3 as the top of the first column in AC but not 1 in the bottom of it.

OpenStudy (phi):

\[\left[\begin{matrix}21 & 5 \\ -10 &1\end{matrix}\right]\]

OpenStudy (phi):

it would be -7 4 for the 1st col of C

OpenStudy (anonymous):

Fair enough, so explain to me the first column of C.

OpenStudy (anonymous):

Does the top have to be the negative double plus or minus 1?

OpenStudy (anonymous):

If the bottom is negative, then plus 1

OpenStudy (anonymous):

If the top is negative then minus 1

OpenStudy (phi):

\[\left[\begin{matrix}1-2b & 7-2c\\ b & c\end{matrix}\right]\]

OpenStudy (anonymous):

But the second column doesn't matter as that is already non zero.

OpenStudy (phi):

just being complete

OpenStudy (anonymous):

Would this be right for the first column? \[\left(\begin{matrix}-a \\-2(a)-1\end{matrix}\right)\]

OpenStudy (anonymous):

Or \[\left(\begin{matrix}-2(-b)+1 \\-b\end{matrix}\right)\]

OpenStudy (anonymous):

\[\left(\begin{matrix}2b+ \\ -b\end{matrix}\right)\]

OpenStudy (anonymous):

\[\left(\begin{matrix}2b+1 \\ -b\end{matrix}\right)\]

OpenStudy (phi):

how are getting these? no.

OpenStudy (anonymous):

If b is 10?

OpenStudy (anonymous):

21 and -10

OpenStudy (anonymous):

\[\left(\begin{matrix}-a \\ 2a-1\end{matrix}\right)\]

OpenStudy (anonymous):

Or in that instance if a is 1 then -1 and 1

OpenStudy (anonymous):

Both examples you gave.

OpenStudy (phi):

for your last -a 2a-1 if a=0 you get 0 -1 and that as your 1st col gives -6 -2 not 3 1

OpenStudy (anonymous):

If the negative is on top then it's multiplying the 3 so the bottom one has to be the negative double minus 1. If the negative is on the bottom it's multiplying the 6 so the bottom has to be the negative double plus 1.

OpenStudy (anonymous):

Only cause a is 0

OpenStudy (anonymous):

In every other instance it works doesn't it?

OpenStudy (anonymous):

-1 1, -2 3, -3 5,

OpenStudy (phi):

think about it some more....I am not leading you astray (though you seem to want to wander off into the woods on this)....

OpenStudy (anonymous):

I'm not saying you are, but my explanation makes more sense to me.

OpenStudy (anonymous):

If the 6 is negative, you need to add one more than double the number of 3s to get it to three.

OpenStudy (anonymous):

Oh I don't care any more. Thank you.

OpenStudy (anonymous):

Thank you for answering, not for making me not care anymore :P